OpenStudy (anonymous):

What is the slope of the following equation? y = x + 5

jimthompson5910 (jim_thompson5910):

The slope of y = mx+b is m We can see from y = x+5 or y = 1x+5 that m = 1 So the slope is 1

OpenStudy (anonymous):

wow thanks, such a simple thing

jimthompson5910 (jim_thompson5910):

np, glad to help out

OpenStudy (anonymous):

What is the slope of the line given by the following equation? y = -5x - 6

jimthompson5910 (jim_thompson5910):

The slope of y = mx+b is m We can see from y = -5x-6 that m = -5 So the slope is -5

OpenStudy (anonymous):

ooo ok i got it now

jimthompson5910 (jim_thompson5910):

yeah, basically look for the number sitting in front of x

OpenStudy (anonymous):

so, if its the Y it would be the second digit?

jimthompson5910 (jim_thompson5910):

not sure what you mean exactly